# Learning High-Order Relationships of Brain Regions

### v2

@@ Abstract
Discovering reliable and informative relationships among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in phenotypic predictions. Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions. We propose that these high-order relationships should be maximally informative and minimally redundant (MIMR). However, identifying such high-order relationships is challenging and under-explored due to the exponential search space and the absence of a tractable objective. In response to this gap, we propose a novel method named HyBRiD which aims to extract MIMR high-order relationships from fMRI data. HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge, which avoids searching in exponential space. HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees. Our comprehensive experiments demonstrate the effectiveness of our model. Our model outperforms the state-of-the-art predictive model by an average of $11.2\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
@@ 1 Introduction
Discovering relations among brain regions toward a specific phenotypic outcome from fMRI signals has been a crucial area in neuroimaging research. Reliable and informative relations help neuroscientists and clinical professionals to better understand brain functions, and thus improve clinical diagnosis and treatments (Kucian et al., 2008, 2006; Li et al., 2015b, a; Satterthwaite et al., 2015; Wang et al., 2016).
However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges. In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019). For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014). Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance. Although recently there have been a few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they are unable to effectively extract meaningful patterns. This is because these methods usually first identify some candidates of high-order relationships and then perform feature selection. The candidates are obtained by enumerating only a small portion of all possible relations, or by clustering methods that are unrelated to the target. As a result, most informative high-order relations are likely excluded at the first stage. This inspires us to solve the problem in an end-to-end manner through a more expressive model.
@@
However, current methods for hypergraph construction, which are mostly based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges. 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023). 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance. Furthermore, although information bottleneck (IB) has been a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs. Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
@@ 3 Related Work
Existing hypergraph construction methods are mostly based on neighbor reconstruction and neighbor distances. For example, the $k$ nearest neighbor-based method (Huang et al., 2009) connects a centroid node and its $k$ nearest neighbors in the feature space to form a hyperedge. Wang et al. (2015); Liu et al. (2017); Jin et al. (2019); Xiao et al. (2019) further refine these neighbor connections through various regularization. However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances. Zhang et al. (2022, 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods. Therefore, they share the same limitations as the aforementioned methods. In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective. Other methods, such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019), are ill-adapted to our context since they require discrete labels. Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model. Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
@@
Information bottleneck (IB) (Tishby et al., 2000) is originally a technique in data compression. The key idea is to extract a summary of data, which contains the most relevant information to the objective. Alemi et al. (2016) first employed an IB view of deep learning. After that, IB has been widely used in deep learning. The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020, 2022; Xu et al., 2021; Wu et al., 2020). Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
@@ 4.1 Learning the Hypergraph by Multi-head Masking
In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and simply aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks. The number of hyperedges $K$ is decided according to the study in Appendix H.1.
@@ 4.2 Optimization Framework
In conclusion, The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7. For how we choose the trade-off coefficient $\beta$, see Appendix H.2 for more discussions.
@@ 5.1 Predictive Performance of Hypereges
where $\boldsymbol{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges. $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the outcome. To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
$r=\operatorname{CPM}([\boldsymbol{w}_{p}\|\boldsymbol{w}_{h}],Y),$ (13)
[…]
@@
Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
Standard pairwise $0.113$ $0.218$ $0.099$ $0.164$ $0.201$ $0.322$ $0.299$ $0.289$
$l_{1}$ hypergraph $0.099$ $0.223$ $0.125$ $0.126$ $0.145$ $0.295$ $0.242$ $0.259$
@@ Results
We report $r$ values by CPM in Table 1 and Table 2. As we can see, on the ABIDE dataset, HyBRiD consistently outperforms all the baselines on different targets, with an average improvement of $8.8\%$ compared to the state-of-the-art model. On the ABCD dataset, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets of ABCD, with an average improvement of $12.1\%$. The results demonstrate our model is able to learn informative hyperedges towards different phenotypic outcomes from fMRI data of various brain states. Rest 1 is the only dataset that our model reaches the second-best. We attribute this imperfection to the discrepancy between validation and test datasets. This is because, through the inspection of the training/validation set performances, we find our model performs better than BrainNetTF on the training and validation dataset. More discussions about it can be found in Appendix F.
@@
HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix I for more runtime details.
@@
[…]
Extra Experiments on Synthetic Dataset
@@ 5.2 Further Analysis
Since CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test. We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
The relationship between hyperedge degree and its significance is shown in Figure  3(b). In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones. It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 3(a).
Comparison with Pairwise Edges
@@
Region Importance
More visualizations of individual hyperedges can be found in Appendix J.
@@ 6 Conclusion
7 Impact Statements
@@ Appendix E Training Details
The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix H.
notation
meaning
value
$lr$
learning rate
$1\times 10^{-3}$
$K$
number of hyperedges
$32$
$\beta$
trade-off coefficients information bottleneck
$0.2$
$[h_{1},h_{2},h_{3}]$
hidden sizes of the dim reduction MLP
$[32,8,1]$
$B$
batch size
$64$
@@ Appendix F Why is HyBRiD not the Best on the Rest 1 Dataset?
Our model reaches the second-best on the Rest 1 dataset. To find out the reason, we inspect the performance on both the training, validation and test datasets. We summarize the performances of our model and the SOTA one in the table below
model train validation test
BrainNetTF $0.967$ $0.484$ $0.334$
HyBRiD (Ours) $0.984$ $0.491$ $0.223$
Table 7: Our model’s performance compared to the SOTA one on the Rest 1 dataset.
We find our model outperforms BrainNetTF on both training and validation datasets but fails on test datasets. Therefore, given that Rest 1 is the most noisy dataset, we argue that the imperfection is because
1) Our model overfits the training dataset more than BrainNetTF.
2) The validation score is higher, indicating the reason is the discrepancy between the training dataset and the test dataset.

### v3

@@ Abstract
Discovering reliable and informative relationships among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in phenotypic predictions. Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions. We propose that these high-order relationships should be maximally informative and minimally redundant (MIMR). However, identifying such high-order relationships is challenging and under-explored due to the exponential search space and the absence of a tractable objective. In response to this gap, we propose a novel method named HyBRiD which aims to extract MIMR high-order relationships from fMRI data. HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge, which avoids searching in exponential space. HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees. Our comprehensive experiments demonstrate the effectiveness of our model. Our model outperforms the state-of-the-art predictive model by an average of $11.2\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections. Source code is available at https://github.com/Graph-and-Geometric-Learning/HyBRiD.
@@ 1 Introduction
Discovering relations among brain regions toward a specific phenotypic outcome from fMRI signals has been a crucial area in neuroimaging research. Reliable and informative relations help neuroscientists and clinical professionals to better understand brain functions, and thus improve clinical diagnosis and treatments (Kucian et al., 2008, 2006; Li et al., 2015b, a; Satterthwaite et al., 2015; Wang et al., 2016). However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges. In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019). For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014). These high-order relationships, cannot always be decomposed into pairwise ones (Battiston et al., 2020, 2021; Bick et al., 2023), thus allowing them to capture information beyond the reach of pairwise ones (Do et al., 2020). Consequently, relying solely on pairwise connectivity, without accounting for the brain’s complex high-order structure, may result in inconsistent findings and low predictive performance across studies. Although recently there have been a few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they are unable to effectively extract meaningful patterns. This is because these methods usually first identify some candidates of high-order relationships and then perform feature selection. The candidates are obtained by enumerating only a small portion of all possible relations, or by clustering methods that are unrelated to the target. As a result, most informative high-order relations are likely excluded at the first stage. This inspires us to solve the problem in an end-to-end manner through a more expressive model and a more appropriate objective.
@@
However, current methods for hypergraph construction, which are mostly based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges. 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023b). 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance. Furthermore, although information bottleneck (IB) has been a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs. Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
@@ 3 Related Work
Existing hypergraph construction methods are mostly based on neighbor reconstruction and neighbor distances. For example, the $k$ nearest neighbor-based method (Huang et al., 2009) connects a centroid node and its $k$ nearest neighbors in the feature space to form a hyperedge. Wang et al. (2015); Liu et al. (2017); Jin et al. (2019); Xiao et al. (2019) further refine these neighbor connections through various regularization. However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances. Cluster-based approaches learn community structure in a graph (Bannadabhavi et al., 2023; Ying et al., 2018), which can also be used to form hyperedges. Zhang et al. (2022, 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods. Therefore, they share the same limitations as the aforementioned methods. In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective. Other methods, such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019), are ill-adapted to our context since they require discrete labels. Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model. Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
@@
Information bottleneck (IB) (Tishby et al., 2000) is a technique in data compression. The key idea is to extract a summary of data, which contains the most relevant information to the objective. Alemi et al. (2016) first employed an IB view of deep learning. After that, IB has been widely used in deep learning. The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020, 2022; Xu et al., 2021; Wu et al., 2020). Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
@@ 4.1 Learning the Hypergraph by Multi-head Masking
In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and simply aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks. The number of hyperedges $K$ is decided according to the study in Appendix I.1.
@@ 4.2 Optimization Framework
In conclusion, The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7. For how we choose the trade-off coefficient $\beta$, see Appendix I.2 for more discussions.
@@ 5.1 Predictive Performance of Hypereges
where $\boldsymbol{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges. $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the outcome.
To evaluate the quality of the learned weights for our model, we replace the pairwise edge weights with the learned high-order weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$, and thus adjust Equation 12 to
$r=\operatorname{CPM}(\boldsymbol{w}_{h},Y),$ (13)
Comparing $r$ to $r^{\prime}$ reflects the quality of learned weights in terms of the prediction performance since it measures the overall correlation between weights and the prediction target. In our model, $\boldsymbol{w}_{h}=\boldsymbol{w}$, which is the learned hyperedge weights.
Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
Standard pairwise $0.113$ $0.218$ $0.099$ $0.164$ $0.201$ $0.322$ $0.299$ $0.289$
$l_{1}$ hypergraph $0.099$ $0.223$ $0.125$ $0.126$ $0.145$ $0.295$ $0.242$ $0.259$
@@
Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
HyBRiDNoMask $0.297_{\pm 0.035}$ $0.274_{\pm 0.057}$ $\underline{0.323_{\pm 0.059}}$ $0.221_{\pm 0.034}$ $0.653_{\pm 0.036}$ $0.599_{\pm 0.059}$ $0.757_{\pm 0.021}$ $0.543_{\pm 0.038}$
HyBRiDRndMask $0.256_{\pm 0.069}$ $0.191_{\pm 0.046}$ $0.255_{\pm 0.080}$ $0.190_{\pm 0.051}$ $0.541_{\pm 0.069}$ $0.514_{\pm 0.038}$ $0.598_{\pm 0.064}$ $0.482_{\pm 0.083}$
HyBRiDSoftMask $\underline{0.343_{\pm 0.042}}$ $\underline{0.314_{\pm 0.040}}$ $0.320_{\pm 0.055}$ $\mathbf{0.245_{\pm 0.061}}$ $\underline{0.707_{\pm 0.042}}$ $\underline{0.662_{\pm 0.058}}$ $\underline{0.796_{\pm 0.031}}$ $\underline{0.655_{\pm 0.030}}$
@@ Results
We report $r$ values by CPM in Table 2 and Table 1. As we can see, on the ABIDE dataset, HyBRiD consistently outperforms all the baselines on different targets, with an average improvement of $8.8\%$ compared to the state-of-the-art model. On the ABCD dataset, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets of ABCD, with an average improvement of $12.1\%$. The results demonstrate our model is able to learn informative hyperedges towards different phenotypic outcomes from fMRI data of various brain states. Rest 1 is the only dataset that our model reaches the second-best. We conduct analyses and propose the potential reasons in Appendix F.
Further, the comparison between our model and the pairwise baseline demonstrates the superiority of incorporating high-order relationships over relying solely on pairwise ones.
@@
HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix J for more runtime details.
@@
We conduct an ablation study on the effect of our masking mechanism. Specifically, we compare our model with $3$ variants: 1) HyBRiDRndMask: Replace the learnable masks with random masks with the same sparsity, initialized at the beginning of training. 2) HyBRiDNoMask: Do not mask at all, which means all nodes and their features are visible to each head. 3) HyBRiDSoftMask: Remove the indicator function and use $p^{k}_{\theta,i}$ directly in Equation 3. Ablation results are shown in Table 3. We find the original HyBRiD and the HyBRiDSoftMask outperform all other variants, which demonstrates the effect of learnable masks. Moreover, the original HyBRiD is better than its soft version HyBRiDSoftMask, which demonstrates our sparse and succinct representations preserve better information than smooth ones. Other ablation studies such as the choices of the number of hyperedges and choices of $\beta$ can be found in Appendix I.2.
Additional Experiments on Synthetic Dataset
@@
Additional Experiments on Model Fit
We further discuss the model’s goodness of fit in Appendix H, with Mean Squared Error (MSE) as the evaluation metric. Our model outperforms the state-of-the-art model in 9 out of 11 datasets, with an average improvement of $11.9\%$.
@@ 5.2 Further Analysis
Since CPM conducts a significance test (details can be found in Appendix B) on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test. We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
The relationship between hyperedge degree and its significance is shown in Figure  4(b). In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones. It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 4(a).
High-order relationships are Better than Pairwise Ones
@@
Hyperedge Case Study
More visualizations about individual region importance can be found in Appendix K.
@@ 6 Conclusion
Additionally, HyBRiD does not offer a method for interpreting complex high-order relationships. This limitation is not specific to HyBRiD, but is a common challenge in analyzing such relationships.We propose a hierarchical strategy that has the potential to interpret them to a certain extent, which is detailed in Appendix L
Impact Statement
@@ Appendix E Training Details
The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix I.
@@
notation meaning value
$lr$ learning rate $1\times 10^{-3}$
$K$ number of hyperedges $32$
$\beta$ trade-off coefficients information bottleneck $0.2$
$[h_{1},h_{2},h_{3}]$ hidden sizes of the dim reduction MLP $[32,8,1]$
$B$ batch size $64$
@@ Appendix F Why is HyBRiD not the Best on the Rest 1 Dataset?
At present, we do not fully understand the reasons for the phenomena observed. The hypotheses outlined below are based on our preliminary observations and analyses.
Hypothesis 1: Sparsity in Resting-State Connectivity
We propose that resting-state connectivity is sparser and more diffuse compared to task-based connectivity. This assertion is refined from our initial claim that the connectivity relevant to the phenotypic outcomes during the resting state is notably sparser than during the task-based state.
Upon examining traditional pairwise connections, we summarized the number of significant connections most related to the phenotypic outcome in Table 7. Our analysis indicates that Rest 1 exhibits the fewest phenotype-related connections, suggesting that these connections are indeed sparse. We did not evaluate high-order relationships due to the lack of ground truth data for such connections. Further investigation is required to validate this hypothesis.
SST $1$ EN-back $1$ MID $1$ Rest $1$ SST $2$ EN-back $2$ MID $2$ Rest $2$
# (connections) $3164$ $3414$ $2989$ $2036$ $3112$ $3014$ $2692$ $3896$
Table 7: The number of significant pairwise edges selected by CPM in different tasks and timepoints.
Hypothesis 2: Comparative Quality of Data Across Timepoints
We hypothesize that the data quality at timepoint 1 is inferior to that at timepoint 2. The performance metrics, as shown in the Table 8, are better at timepoint 2 than at timepoint 1. Moreover, timepoint 1 exhibits more severe overfitting compared to timepoint 2. This discrepancy may be attributed to the higher relevance of the information (pertaining to the outcome) and lower noise levels at timepoint 2. This hypothesis also requires further verification through detailed investigations and experiments.
SST $1$ EN-back $1$ MID $1$ Rest $1$ SST $2$ EN-back $2$ MID $2$ Rest $2$
train $0.988$ $0.984$ $0.984$ $0.984$ $0.982$ $0.977$ $0.978$ $0.976$
test $0.361$ $0.348$ $0.386$ $0.223$ $0.738$ $0.714$ $0.816$ $0.730$
Table 8: Performances on training and test dataset in different tasks and timepoints.
Conclusion
Brain activities during resting states are not driven by external tasks, leading to more diffuse and less predictable patterns of activation. The low data quality of ABCD timepoint 1 (compared to timepoint 2) even intensifies this issue. Therefore, high-order relations might not be a good inductive bias on the Rest 1 dataset since the connections might be much more sparse and involve fewer nodes.